# VoCoT: Unleashing Visually Grounded Multi-Step Reasoning in Large Multi-Modal Models

### v2

@@ 3.2 Training
The training of VolCano undergoes three stages summarized below. For more details on the utilized and the implementation settings, please refer to Section 4.1 and Appendix A.3.
@@
The first stage aims to align visual representations with the LLM backbone. We utilize image-caption pairs from LLaVA-Pretrain [31; 29]. Both the visual encoder and the LLM backbone are frozen while the parameters in the connection module are updated.
@@
Following the alignment stage, the model is trained to adapt to multi-modal interleaved sequences and visually grounded object representations with RefBind. Three types of sequences are considered: (1) Image-caption pair constitutes the simplest form of multi-modal sequences. Different from the first stage, we utilize captions collected by ALLaVA [4] which provide detailed descriptions. (2) Multi-modal document includes multiple image-text pairs in a sequence. Based on the relevance between images and text, we filter a subset of documents from MMC4 [63]. (3) Grounded image caption further annotates the coordinates of objects in the caption. The representations of included objects are extended to the visually grounded format consistent with VoCoT. Flickr30K Entities [39] and a filtered subset of GRIT [38] are adopted. Besides the connection module, the language backbone is trained to model the multi-modal sequences.
@@
Based on the pre-trained model, the model is fine-tuned to follow instructions in multi-modal context and perform multi-step reasoning with VoCoT. We supplement the constructed VoCoT-Instruct-80K introduced in Section 2.2 and referring expression data [18; 6] to existing visual instruction data [29], which does not include CoT-form data. Similar to the last stage, both the language backbone and the connection module are fine-tuned.
Training Objective
The training objective of VolCano is unified as a causal modeling loss:
$\mathcal{L(\theta)}=-\mathbb{E}_{x_{i}\sim D}\Big{[}\frac{1}{|v_{i}|}\sum_{t\in v_{i}}\log P_{\theta}(x_{t}|\{x_{j}\}_{j<t})\Big{]}$ (1)
where $x_{i}$ is the constructed sequence in $D$, the training dataset. $P_{\theta}$ is the probability distribution modeled by the model and $\theta$ denotes the parameters. For each sequence $x_{i}$, $v_{i}$ indicates the set of indexes for tokens that need to be predicted. In practice, $v_{i}$ only contains textual tokens, including special tokens, coordinates, and natural text. In terms of pre-training data, all textual tokens are included in $v_{i}$, whereas for instruction-tuning data, only tokens in the responses are considered in $v_{i}$.
@@ 4.1 Experiment Settings
In this paper, we build VolCano with the pre-train ViT-L/14 CLIP [40] visual encoder and Mistral-7B [16] as the language backbone. Compared to Vicuna [8] and LLaMA [45], Mistral offers superior reasoning abilities in text, making it a better choice for modeling VoCoT. The input resolution of images are set to $336\times 336$. Experiments are conducted using 8 NVIDIA A100 GPUs. The model is trained for one epoch in each stage, the complete training process takes approximately 90 hours. Detailed parameter settings are provided in Appendix A.3.
Tasks and Datasets
To validate the effectiveness and versatility of the multiple-reasoning paradigm formulated by VoCoT, we adopt different tasks across various scenarios for assessment:
  • General VQA comprehensively evaluates various capabilities of LMMs by integrating different types of instructions and questions. We consider the commonly adopted GQA [15], MMBench [33], and SEED [21] benchmarks. For SEED, we observe that different evaluation methods are used by different models. To conduct a fair comparison, we re-evaluate the compared models with the likelihood evaluation method described in the original paper [21].
  • Visual spatial reasoning requires models to distinguish the spatial relationship between objects. VSR [28] and EmbSpatial [10] are utilized for assessment.
  • Hallucination tasks. Since hallucination significantly hinders effective reasoning. We utilize POPE [24] and AMBER [48] to diagnose the reliability of the generated outputs.
  • Compositional tasks in multi-modal problems require solution logic that includes multi-step analysis and reasoning. CLEVR [17], V-Star [53], and Winoground [44] are applied. Following Reform-Eval [25], we formulate Winoground as a binary classification task of selecting the correct description from two similar captions based on an image.
Except for AMBER, where CHAIR [48] is the evaluation metric, accuracy is reported for other datasets. For details on the evaluation datasets and evaluation methods, please refer to Appendix A.4.
Baseline Models
We compare VolCano with existing LMMs based on single-step reasoning with ~7B parameters, including BLIP-2 [22], InstructBLIP [9], Shikra [6], mPLUG-Owl2 [59], MiniGPT-v2 [5], Qwen-VL-Chat [2], and LLaVA-1.5 [29]. Furthermore, based on our framework and data with VoCoT-Instruct-80K removed, we construct a single-step reasoning baseline model, VolCano-SE, for a strict comparison. The models listed before, including VolCano, take a single image as input. Additionally, we include another series of SOTA LMMs that enhance input resolution by splitting a single image into multiple sub-images: LLaVA-1.6 [30], Deepseek-VL [34], Monkey [27].
Notations
@@
4.2 Main Results
Table 2: Comparison with SOTA LMMs on 10 benchmarks.
Model General VQA Spatial Reasoning Compositional Tasks Hallucination
@@ 4.2 Main Results
Models with single-image inputs
BLIP-2-7B $224^{2}$ 1.3B 41.00 - - 50.90 29.50 - 37.70 - - -
InstructBLIP-7B $224^{2}$ 1.3B 49.20 36.00 - 52.10 29.00 - 34.02 - 72.10 8.80
Shikra-7B $224^{2}$ 0.3B - 58.80 - - 32.60 - - - 83.10 -
mPLUG-Owl2-7B $448^{2}$ 0.3B 56.10 64.50 59.99 - 36.72 43.22 36.12 63.38 - 10.60
MiniGPT-v2-7B $448^{2}$ 1.3B 60.10 55.14 51.50 62.90 45.10 46.23 33.19 62.00 80.50 -
Qwen-VL-Chat $448^{2}$ 1.9B 57.50 60.60 64.70 - 38.68 53.20 45.80 - 84.70 5.50
LLaVA1.5-7B $336^{2}$ 0.3B 62.00 64.30 53.80 64.24 42.43 43.73 48.68 55.31 84.50 7.80
Models with multiple-image inputs
GPT-4V $2048^{2}$ 75.80 71.60 68.24 26.65 51.90 55.00 83.75 82.00 4.60
Table 2 presents a thorough evaluation of existing LMMs. Several insights can be gleaned: (1) Among models with single-image inputs, VolCano achieves the best performance in most datasets except SEED-Image, with the advantage being more pronounced in complex tasks. (2) Introducing VoCoT in LMMs effectively alleviate the hallucination issue. Section 4.3 delves deeper into how VoCoT contributes to reliable and visually grounded reasoning. (3) Comparing VolCano with LMMs that enhance input resolution by introducing multiple-image inputs, we observe that these methods primarily improve the performance in general VQA and V-Star, as V-Star provides high-resolution input images [53]. However, in tasks that require complex reasoning, the improvement brought by higher resolutions becomes less significant. VolCano either exceeds or perform comparably with these models in such tasks, indicating the superiority of introducing multi-step reasoning over enriching the input information in these scenarios. We leave exploring combining these two vertical research directions—enhancing input resolution and introducing multi-step reasoning–as future work. Overall, the experimental results validate the effectiveness of VoCoT-based multi-step reasoning in various scenarios, particularly in tasks demanding complex reasoning.
4.3 Comparison between CoTs in Different Formats
In this section, we demonstrate the effectiveness of the VoCoT format by comparing it with different CoT formats: (1) Zero-Shot CoT directly prompts VolCano-SE model to include a object-centric reasoning path without training; whereas (2) Text CoT represents objects with only text descriptions; (3) Coor. CoT augments Text CoT with coordinates; and (4) Sub-Img CoT encodes sub-images as visual representations of objects rather than using RefBind, during the training process.
Table 3 lists the results. Firstly, the zero-shot multi-step reasoning capability of VolCano-SE is limited and likely to exhibit hallucinations, highlighting the necessity to construct visual CoT tuning data. Secondly, CoT expressed only in texts is also affected by hallucinations, handling spatial reasoning well where each type of object appears only once, but failing to manage more complex scenarios. Thirdly, introducing coordinates grounds the thoughts to visual signals, mitigating hallucination and improving performance across various tasks. Furthermore, the representations obtained by RefBind effectively help the model utilize visual signals of objects. However, simply encoding sub-images may backfire by including redundant information. Overall, VoCoT is the most suitable format.
Table 3: Comparison between different CoT formats. The “Obj-Format” column indicates the representation format of objects. T, C, S, and R are short for texts, coordinates, sub-images, RefBind.
@@ 4.3 Comparison between CoTs in Different Formats
4.4 Ablation on Constructed Datasets
In Table 4, we explore the role of three types of data in VoCoT-Instruct-80K. The results implies that: (1) Type 1, the GQA-based data, is precise but limited in terms of diversity. Models trained solely on Type 1 data produces the fewest hallucinations but struggle to handle diverse questions. (2) Type 2 and Type 3 data effectively help the model generalize across various instructions. Nevertheless, removing Type 1 data from the training set increases the risk of hallucinations. (3) Introducing multi-modal interleaved data in Stage 2 significantly improve the performance. In summary, three types of data should be jointly utilized, and interleaved data need to be incorporated for pre-training.
@@ 4.5 Further Analysis
Figure 4 compares the performance of VolCano-SE and VolCano on questions of varying difficulty in CLEVR [32]. Based on the fitted curves and confidence intervals, it is evident that as the number of required reasoning steps increases, the advantage of multi-step reasoning becomes more pronounced.
Grounding Capabilities of VolCano
Table 5: Referring expression results.
Model RefCOCOg CLEVR Ref
Method Res. val test
Shikra-7B $224^{2}$ 82.3 82.2 -
MiniGPT-v2-7B $448^{2}$ 84.2 84.3 24.9
Qwen-VL-Chat-7B $448^{2}$ 85.9 86.3 22.4
LLaVA1.5-7B $336^{2}$ 70.5 71.0 6.7
VolCano-7B $336^{2}$ 83.1 84.4 34.0
In this section, we attempt to separately analyze two aspects of capabilities in VolCano: visual grounding and reasoning. Referring expression task in RefCOCOg [18] and CLEVR-Ref [32] is adopted to evaluate the former capability. According to Table 5, VolCano with lower input resolution demonstrates comparative performance on RefCOCOg which is an in-domain dataset. In the out-domain and complex dataset, CLEVR-Ref, VolCano significantly outperforms other models. These results indicate that VolCano possesses strong and generalizable grounding ability.
Reasoning Capabilities of VolCano
Regarding the reasoning process, we divide it into two sub-processes: analysis and judgment, where the former constructs reasoning paths and the latter provides conclusions. We conduct experiments on VSR: VolCano serves as the analyzer to provide reasoning paths based on which the conclusions are drawn by different judgers. In VSR, there is one-to-one correspondence between objects in the image and textual descriptions, LLMs can effectively make judgments based on the textual reasoning paths generated by VolCano after removing the coordinates and visual information. In addition, we construct VolCanoV, a Vicuna-based variant of VolCano, to investigate the impact of language backbones. According to Table 6, a more powerful judger can more effectively utilize the reasoning paths generated by Vocalno. When using VolCano as the analyzer and GPT-4 as the judger, the performance even surpasses that of GPT-4V. This indicates that VoCalno possesses the capability to provide effective reasoning paths, but its judgment ability is limited. Furthermore, it can be observed that the judgment ability of VolCano is mainly limited by the backbone. Generally, the experiments highlight the importance of the language backbone and reveal the potential of further applying VoCoT on stronger LLM backbones.
Table 6: Performance with different analyzers and judgers on VSR dataset.
Analyzer VolCanoV VolCano GPT-4V
Accuracy 63.5 64.5 67.2 67.2 65.1 67.8 73.8 68.24
@@
As shown in Figure 5, in AMBER, VolCano provides a visually grounded description with no hallucinations. In CLEVR, VoCalno can infer effective multi-step reasoning paths towards the answer. Please refer to Appendix B.1 for more cases from different datasets.
@@ 5.1 Large Multi-Modal Models
Architecture of Large Multi-Modal Models
A vast amount of research has emerged, focusing on adapting large language models (LLMs) to handle multi-modal tasks. Initially, researchers treated large language models (LLMs) as intelligent agents capable of using various tools. They trained these models to invoke fundamental vision models, enabling them to complete multi-modal tasks such as object detection, image description, and visual question answering [52; 43; 56]. More recent works have focused on directly aligning the visual and textual modalities. To achieve this, researchers design various connection modules, including linear projections [31; 11; 5; 30; 29; 34; 6; 26], Q-Former [22; 14; 62; 9], and cross-attention-based re-sampler (abstractor) [1; 58; 2; 27; 59].
@@
To enhance the instruction-following abilities of large language models (LLMs) in multi-modal contexts, a significant line of research has focused on converting existing task-oriented datasets into instructional data [9; 29; 23]. Utilizing models akin to GPT-4, such as MiniGPT4 [62], researchers bootstrap high-quality data from the model outputs, and LLaVA [30] transforms image-caption pairs into dialogue data. Further efforts are directed towards improving both the quality and scale of the generated data [4; 61; 47].
5.2 Multi-Step Reasoning within Large Pre-trained Models
CoT in Large Language Models
Chain of Thoughts (CoT) is a series of prompting techniques designed to facilitate large language models (LLMs) in addressing complex problems by guiding them through intermediate steps. CoT is first proposed through in-context learning [51], followed by the introduction of zero-shot CoT [19], Auto-CoT [60] and self-consistency mechanism [50]. Subsequently, CoT are adapted to consider complex formats [57; 3; 7].

### v3

@@ 3.2 Training
The training of VolCano undergoes three stages:
@@
The first stage aims to align visual representations with the LLM backbone. We utilize image-caption pairs from LLaVA-Pretrain Liu et al. (2023b). Only the parameters in the connection module are updated.
@@
Following the alignment stage, the model is trained to adapt to multi-modal interleaved sequences and visually grounded object representations with RefBind. Three types of sequences are considered: (1) Image-caption pair constitutes the simplest form of multi-modal sequences. We utilize ALLaVA-Caption Chen et al. (2024) which provides detailed descriptions. (2) Multi-modal document includes multiple image-text pairs in a sequence. Based on the relevance between images and text, we filter a subset of documents from MMC4 Zhu et al. (2024). (3) Grounded image caption further annotates the coordinates of objects in the caption. We extend the object representations to the visually grounded format consistent with VoCoT. Flickr30K Entities Plummer et al. (2015) and a subset of GRIT Peng et al. (2023) are adopted. Both the connection module and the LLM backbone are trained to model the multi-modal sequences.
@@
The pre-trained model is further fine-tuned to follow instructions in multi-modal contexts and perform multi-step reasoning with VoCoT. We supplement the constructed VoCoT-Instruct-80K and referring expression data Kazemzadeh et al. (2014); Chen et al. (2023b) to the existing non-CoT-form visual instruction data Liu et al. (2023b). We update the LLM backbone and connection module in this stage.
Table 1 summarizes the data mixtures used during the three training stages of VolCano. For more details, please refer to Section 4.1 and Appendix A.3.
@@ 4 Experiments
Stages Data Type Source Size
Stage 1 Image-Caption LLaVA 558k
Stage 2 Image-Caption ALLaVA 695k
Grounded Image-Caption GRIT 756k
Flickr30k 148k
Multimodal Document MMC4 890k
Stage 3 Visual Instruction LLaVA 612k
Referring Expression Shikra-RD 6k
RefCOCO 42k×$\times$3
g-RefCOCO 79k
VoCoT This Work 80k
Table 1: The training data mixtures used by VolCano.
@@ 4.1 Experiment Settings
We build VolCano with the pre-trained ViT-L/14 CLIP Radford et al. (2021) visual encoder and Mistral-7B Jiang et al. (2023) as the baseline backbone. In addition, we explore the impact of a more powerful LLM backbone in our framework, constructing VolCanoQ2 based on Qwen2-7B Yang et al. (2024). Detailed parameter settings are provided in Appendix A.3. To save resources, we merely evaluate VolCanoQ2 in the main experiments and primarily focus on the Mistral-based VolCano in further analysis.
Evaluation Benchmarks
To validate the effectiveness and versatility of the VoCoT framework, we adopt different tasks across various scenarios for assessment: (1) General VQA benchmarks, including GQA Hudson and Manning (2019), MMBench Liu et al. (2023c), and SEED Li et al. (2023a); (2) Composite tasks requiring multi-step analysis and composite capabilities, such as visual spatial reasoning in VSR Liu et al. (2023a) and EmbSpatial Du et al. (2024), visual search in V-Star Wu and Xie (2023), complex reasoning in CLEVR Johnson et al. (2017) and Winoground Thrush et al. (2022), and complex referring expression in CLEVR-Ref Liu et al. (2019); (3) Hallucination benchmarks, including POPE Li et al. (2023d) and AMBER Wang et al. (2023b), to evaluate whether VoCoT can mitigate hallucinations. AMBER uses CHAIR Wang et al. (2023b) as the evaluation metric while accuracy is reported for other datasets. Details on the evaluation processes are provided in Appendix A.4.
Baselines
We compare VolCano with existing LMMs with ~7B parameters, as listed in Appendix A.5. For strict comparison, we construct a baseline model, VolCano-SE, which is based on the same architecture as VolCano but without VoCoT-Instruct-80K training data, so it can only perform single-step reasoning. We divide models into two groups for comparison: models based on baseline backbones (LLaMA-1,2, Vicuna, Qwen, and Mistral) and models based on advanced backbones (LLaMA-3 and Qwen2). We focus on models with single-image inputs in the main part. Please refer to Appendix B.1 for comparison and discussion involving models that use multiple additional sub-images as inputs for resolution enhancement.
Model General VQA Composite Tasks Hallucination
@@
Models based on baseline LLM backbones
InstructBLIP-7B $224^{2}$ 1.3B 49.20 36.00 - 52.10 33.41 - 34.02 - - 72.10 8.80
Shikra-7B $224^{2}$ 0.3B - 58.80 - - 34.75 - - - - 83.10 -
mPLUG-Owl2-7B $448^{2}$ 0.3B 56.10 64.50 59.99 - 36.72 43.22 36.12 63.38 - - 10.60
MiniGPT-v2-7B $448^{2}$ 1.3B 60.10 55.14 51.50 62.90 43.85 46.23 33.19 62.00 24.90 80.50 -
Qwen-VL-Chat $448^{2}$ 1.9B 57.50 60.60 64.70 - 38.68 53.20 45.80 - 22.35 84.70 5.50
LLaVA1.5-7B $336^{2}$ 0.3B 62.00 64.30 53.80 64.24 42.43 43.73 48.68 55.31 6.70 84.50 7.80
VILA-7B $336^{2}$ 0.3B 62.30 61.50 60.40 66.02 38.05 47.60 46.22 66.37 - 84.50 10.50
VisCoT-7B $336^{2}$ 0.3B 63.00 63.82 63.23 - 37.01 53.15 61.76 56.40 32.05 86.10 7.20
VolCano-SE $336^{2}$ 0.3B 59.91 61.15 54.15 63.42 36.14 51.70 44.96 64.00 21.70 84.50 6.70
VolCano $336^{2}$ 0.3B 64.40 68.10 64.50 67.18 58.29 56.17 58.40 68.37 33.95 86.50 4.60
Models based on advanced LLM backbones
VILA1.5-8B $384^{2}$ 0.4B 63.50 64.38 64.41 53.76 54.95 55.22 58.74 66.00 - 84.90 8.50
Bunny-8B V1.0 $384^{2}$ 0.4B 64.00 70.86 67.59 65.71 53.54 54.47 58.32 68.50 - 86.40 7.40
VolCano-SEQ2 $336^{2}$ 0.3B 62.23 66.87 64.51 69.37 55.19 51.58 56.30 66.63 23.90 85.20 8.00
VolCanoQ2 $336^{2}$ 0.3B 64.60 71.61 66.95 74.22 59.86 56.78 62.81 68.78 34.00 86.60 4.40
GPT-4V $2048^{2}$ 75.80 71.60 68.24 36.07 51.90 55.00 83.75 - 82.00 4.60
@@
Table 3: Comparison between different CoT formats. The “Obj-Format” column indicates the representation format of objects. T, C, S, and R are short for texts, coordinates, sub-images, and RefBind representations.
4.2 Main Results
Table 2 presents a thorough evaluation of existing LMMs. Several insights can be gleaned: (1) By comparing VolCano and VolCano-SE, it demonstrates that VoCoT effectively mitigates hallucinations and brings consistent improvement across all benchmarks. Section 4.3 delves deeper into how VoCoT contributes to reliable and visually grounded reasoning. (2) Across different datasets, VolCano and VolCanoQ2 achieve the best or second-best results within their respective group, where the advantages are more pronounced in composite tasks. On benchmarks like CLEVR and EmbSpatial, VolCano with a limited scale even outperforms powerful GPT-4V. (3) Furthermore, we compare two multi-modal CoT methods: VisCoT and VoCoT. VisCoT Shao et al. (2024) designs a simple two-step reasoning process: first searching for a single relevant region and then answering based on the detected region. Experimental results imply that VisCoT merely performs better on V-Star because the questions in V-Star perfectly align with the two-step search ligic of VisCoT. However, VisCoT falls short in other complex scenarios that involve interaction between multiple objects, indicating that VoCoT is a more generalizable format of multi-modal CoT. Overall, the experimental results validate the effectiveness of VoCoT-based multi-step reasoning in various scenarios. In addition, we find that VoCoT could seamlessly generalize to other scenarios including scene-text-centric tasks, please refer to the results and analysis in Appendix B.2.
Figure 4: Qualitative analysis to compare VoCoT and text-only CoT. Hallucinations are underlined.
@@ 4.2 Main Results
Analyzer VolCanoV VolCano VolCanoQ2 GPT-4V
Accuracy(%) 63.5 64.5 67.2 67.2 65.1 67.8 73.8 74.2 68.2
Table 5: Performance with different analyzers and judgers on the VSR benchmark.
4.3 Comparing CoTs in Different Formats
We validate the effectiveness of the VoCoT format by comparing it with different CoT formats: (1) Zero-Shot CoT directly prompts VolCano-SE to think step-by-step without training; (2) Text CoT represents objects with only text descriptions; (3) Coor. CoT augments Text CoT with coordinates; and (4) Sub-Img CoT encodes sub-images as representations of objects rather than using RefBind.
Table 3 lists the results. Firstly, the zero-shot multi-step reasoning capability of VolCano-SE is limited. It is likely to exhibit hallucinations, highlighting the necessity to construct visual CoT tuning data. Secondly, CoT expressed only in texts is also affected by hallucinations, handling spatial reasoning well where each type of object appears only once, but failing to manage more complex scenarios. Thirdly, introducing coordinates grounds the thoughts to visual signals, mitigating hallucination and improving performance across various tasks. Furthermore, representations obtained by RefBind effectively help the model to utilize visual signals of objects. In contrast, the performance of Sub-Img CoT is overall inferior to that of Coor. CoT, which supports our claim in Section 3.1: simply encoding each object as an sub-image may introduce redundant information and degrade the performance.
Besides quantitative results, we present cases in Figure 4. We observe that text-only CoT is limited in terms of: (i) It may fail to accurately find/locate target object (Case 1). (ii) It is unable to leverage object-level visual information for inferring object attributes, as VoCoT does through RefBind (Case 3). (iii) It cannot resolve ambiguity between multiple objects (Case 4). (iv) Lack of interpretability. In general, it is crucial to ground the reasoning process to the visual information and VoCoT is the most suitable format.
4.4 Ablation on the Constructed Dataset
In Table 4, we explore the role of three types of data in VoCoT-Instruct-80K. The results implies that: (1) Type 1, the GQA-based data, is precise but limited in terms of diversity. Models trained solely on Type 1 data produces the fewest hallucinations but struggle to handle diverse questions. (2) Type 2 and 3 data effectively help the model generalize across various instructions. Nevertheless, totally removing Type 1 data will increase the risk of hallucinations. (3) Introducing multi-modal interleaved data in Stage 2 leads to a significant improvement. In summary, interleaved pre-training data and three types of VoCoT data should be jointly utilized.
@@ 4.5 Further Analysis
Figure 5 compares the performance of VolCano-SE and VolCano on questions of varying difficulty in CLEVR. The fitted curves and confidence intervals imply that as the number of required reasoning steps increases, the advantage of multi-step reasoning becomes more pronounced.
@@
Disentangling Multi-Modal Reasoning
Our preliminary study finds that sometimes VolCano generates reasonable reasoning paths but fail to infer the correct answer. Therefore, we split the reasoning process into two sub-processes: analysis and judgment, where the former constructs reasoning paths and the latter provides conclusions. We conduct experiments to combine different analyzers and judgers on the VSR benchmark, where each object category corresponds to a single object in the image, allowing us to use text-only LLMs to judge based on the paths analyzed by VolCano. In Table 5, VolCanoV represents the Vicuna-based VolCano to explore the impact of LLM backbones. The results indicate that the judger plays a important role. The path analyzed by VolCano help GPT-4 make better decisions (73.8%) than GPT-4V (68.2%). However, the overall capability of VolCano is upper-bounded by the judgement ability of its LLM backbone. Comparison between VolCanoV, VolCano, and VolCanoQ2 further reveals the potential of applying VoCoT on stronger LLM backbones.
@@
Examples in Figure 6 show that VolCano provides a visually grounded description with no hallucinations in AMBER. In CLEVR, VolCano infers effective reasoning paths towards the answer. See Appendix B.8 for more cases in other datasets.
@@ 5.1 Large Multi-Modal Models
Architecture of LMMs
A vast amount of research has emerged, focusing on adapting LLMs to handle multi-modal tasks. Initially, researchers treat LLMs as intelligent agents capable of using various tools. They train or prompt LLMs to invoke fundamental vision models, enabling them to complete multi-modal tasks such as captioning and VQA Wu et al. (2023); Yang et al. (2023c). Recent methods directly align the visual and textual representations in a unified backbone. To achieve this, various connection modules are designed, including MLP Liu et al. (2024b, 2023b), Q-Former Li et al. (2023b); Dai et al. (2023), and cross-attention layers Alayrac et al. (2022); Wang et al. (2023c).
@@
To enable LMMs to follow instructions in multi-modal contexts, a line of research has focused on converting existing academic datasets into instruction data Dai et al. (2023); Li et al. (2023c). In addition, researchers also use powerful tools like GPT-4V to assist in bootstrapping and generating high-quality data Zhu et al. (2023); Liu et al. (2024b). Further efforts are devoted towards improving both the quality and scale of the generated data Chen et al. (2024); Zhao et al. (2023); Wang et al. (2023a).
Visual Grounding in LMMs
Beyond text descriptions, some researchers further empower LMMs to understand and extract fine-grained visual information through visually grounded representations like coordinates Peng et al. (2023); Chen et al. (2023b), masks Yuan et al. (2024) and visual prompts Cai et al. (2024). However, most of these LMMs still rely on single-step and text-only reasoning, failing to effectively leverage fine-grained information during the reasoning process.
5.2 Multi-Step Reasoning
CoT in LLMs
Chain of Thoughts (CoT) is a series of prompting techniques designed to facilitate LLMs in addressing complex problems by guiding them through intermediate steps. CoT is first proposed through in-context learning Wei et al. (2022), followed by the introduction of zero-shot CoT Kojima et al. (2022), Auto-CoT Zhang et al. (2022) and self-consistency Wang et al. (2023d). Subsequently, CoT are extended to more complex formats Yao et al. (2024); Besta et al. (2024).